# Finite-temperature quantum topological order in three dimensions

### v1

@@ Abstract
We identify a three-dimensional system that exhibits long-range entanglement at sufficiently small but nonzero temperature—it therefore constitutes a quantum topological order at finite temperature. The model of interest is known as the fermionic toric code, a variant of the usual 3D toric code, which admits emergent fermionic point-like excitations. The fermionic toric code, importantly, possesses an anomalous 2-form symmetry, associated with the space-like Wilson loops of the fermionic excitations. We argue that it is this symmetry that imbues low-temperature thermal states with a novel topological order and long-range entanglement. Based on the current classification of three-dimensional topological orders, we expect that the low-temperature thermal states of the fermionic toric code belong to an equilibrium phase of matter that only exists at nonzero temperatures. We conjecture that further examples of topological orders at nonzero temperatures are given by discrete gauge theories with anomalous 2-form symmetries. Our work therefore opens the door to studying quantum topological order at nonzero temperature in physically realistic dimensions.
@@
The Hamiltonian terms are mutually commuting and thus define a Pauli stabilizer code. They further satisfy the following relations on a torus:
@@
$\|\rho_{\rm{f\hskip 0.35002ptTC}}-\sigma_{\rm SRE}\|_{1}=1-O(L^{-\infty}).$ (5)
@@
Step one: The first step is to define a QLC that approximately maps the thermal state $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to the loop-less state $\rho_{\varnothing}$. We accomplish this with a QLC $\mathcal{C}$, composed of five layers of QLCs. In each layer, we clean up loop excitations on disjoint patches of linear size at least $\ln(L)^{2}$. After applying all five layers, the only loop excitations that are left behind must have lengths larger than $\ln(L)^{2}$. These large loop excitations are exponentially suppressed in the thermal state at low temperatures. Thus, the QLC approximately maps $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to $\rho_{\varnothing}$, in terms of the fidelity of mixed states. We note that, alternatively, a QLC based on the sweep decoder of Ref. Kubica and Preskill (2019) can be used to remove the small loop excitations, as elaborated on in the Supplemental Materials (SM).
@@
Note that the temperature $T_{0}\approx 1.24$ is below the critical temperature $T_{c}\approx 1.31$ found in Ref. Castelnovo and Chamon (2008).
@@
Since $\mathcal{C}$ is a QLC, the state $\mathcal{C}(\sigma_{\text{SRE}})$ is also SRE, up to a tensor product with the identity, as spelled out in the SM. In particular, there exists a SRE state of the form $\tilde{\sigma}_{\text{SRE}}=\mathcal{C}(\sigma_{\text{SRE}})\otimes I$. Taking a tensor product with the identity does not change the fidelity, so we may write:

### v2

@@ Abstract
We identify a three-dimensional system that exhibits long-range entanglement at sufficiently small but nonzero temperature—it therefore constitutes a quantum topological order at finite temperature. The model of interest is known as the fermionic toric code, a variant of the usual 3D toric code, which admits emergent fermionic point-like excitations. The fermionic toric code, importantly, possesses an anomalous 2-form symmetry, associated with the space-like Wilson loops of the fermionic excitations. We argue that it is this symmetry that imbues low-temperature thermal states with a novel topological order and long-range entanglement. Based on the current classification of three-dimensional topological orders, we expect that the low-temperature thermal states of the fermionic toric code belong to a phase of matter that, in the context of equilibrium phases of matter, only exists at nonzero temperatures. We conjecture that further examples of topological orders at nonzero temperatures are given by discrete gauge theories with anomalous 2-form symmetries. Our work therefore opens the door to studying quantum topological order at nonzero temperature in physically realistic dimensions.
@@
The Hamiltonian terms are mutually commuting and thus define a Pauli stabilizer code Gottesman ; Nielsen and Chuang (2010). They further satisfy the following relations on a torus:
@@
$\frac{1}{2}\|\rho_{\rm{f\hskip 0.35002ptTC}}-\sigma_{\rm SRE}\|_{1}=1-O(L^{-\infty}).$ (5)
@@
Step one: The first step is to define a QLC that approximately maps the thermal state $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to the loop-less state $\rho_{\varnothing}$. We accomplish this with a QLC $\mathcal{C}$, composed of five layers of QLCs. In each layer, we clean up loop excitations on disjoint patches of linear size at least $\ln(L)^{2}$. After applying all five layers, the only loop excitations that are left behind must have lengths larger than $\ln(L)^{2}$. These large loop excitations are exponentially suppressed in the thermal state at low temperatures. Thus, the QLC approximately maps $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to $\rho_{\varnothing}$, in terms of the fidelity of mixed states. We note that, alternatively, a QLC based on the sweep decoder of Ref. Kubica and Preskill (2019) can be used to remove the small loop excitations, as elaborated on in the Supplemental Material (SM).
@@
We note that the finite-temperature phase transition of the fTC occurs at the critical temperature $T_{c}\approx 1.31$. This follows from adapting the arguments of Refs. Castelnovo and Chamon (2008); Weinstein et al. (2020) to the fTC, which also imply that the transition belongs to the 3D Ising universality class. The temperature $T_{0}\approx 1.24$ is therefore below the critical temperature. However, we expect that the LRE persists up to $T_{c}$, and that the value of $T_{0}$ is a consequence of our approximations and a suboptimal algorithm for cleaning up loops.
@@
Since $\mathcal{C}$ is a QLC, the state $\mathcal{C}(\sigma_{\text{SRE}})$ is also SRE, up to a tensor product with the identity, as spelled out in the SM. In particular, there exists a SRE state of the form $\tilde{\sigma}_{\text{SRE}}=\frac{1}{2^{|A|}}\mathcal{C}(\sigma_{\text{SRE}})\otimes I_{A}$, where $|A|$ is the number of ancillary qubits. Taking a tensor product with the identity does not change the fidelity, so we may write: